# Fundamental examples of Reeb spaces of smooth functions defined from two graphs of smooth functions with same asymptotic behaviors

### v2

@@ Abstract.
Hereafter, we also abuse some notation from Proposition 1 and its proof. We choose a point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$. Let $p_{D}:=(p_{D,1},p_{D,2})$. We consider the point $p_{D,c_{1},q}:=(p_{D,1},q_{D,1})$, where $q_{D,1}$ is chosen to be a point with $c_{1}(q_{D,1})=p_{D,1}\leq p_{D,2}$ and can be uniquely defined. The set $\{(p_{D,1},t)\mid q_{D,1}\leq t\leq p_{D,2}\}$ is in $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$. We can consider some path from $p_{D,c_{1},q}$ to another point $p_{s,c_{1},q}$ in $\{(x_{1},x_{2})\mid x_{2}<b_{{\rm m},0},c_{1}(x_{2})=x_{1}\}$ and have the path departing from $p_{D}$ obtained by joining the set $\{(p_{D,1},t)\mid q_{D,1}\leq t\leq p_{D,2}\}$, which is a one-point set or a segment, and the path in $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$ in such a way that the restriction of ${\pi}_{2,1}$ to the latter path is injective and that the point $(p_{s,c_{1},q},{(0)}_{j=1}^{m-1})\in{{\mathbb{R}}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$.
We discuss a fact obtained from the argument on the assumption (1) of Proposition 3 here. First, the assumption (2) of Proposition 3 is essentially same as (1) of Proposition 3. From the assumptions (1, 2) of the present theorem, for a (more) general point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\overline{{D}_{m,c_{1},c_{2}}}$, we can have an oriented path from $p_{e}$ to a vertex of the Reeb graph the orientation at each point of which is same as the orientation induced from the Reeb digraph, as we first argue in the proof of Proposition 1 (and Proposition 2). The assumptions (3, 4) of the present theorem are regarded to be a kind of duals to the assumptions (1, 2) of the present theorem. We can argue as in the proofs of Propositions 1 and 2 to complete the proof of the sufficiency.
@@
Theorem 3 is also regarded as an extension of Proposition 1, where Propositions 2 and 3 with their proofs being important. We only show an example for Proposition 3.
@@
The case $c_{1}(x):=\frac{t_{0}}{x^{2}+1}$ ($0<t_{0}<1$) and $c_{2}(x):=\frac{1}{x^{2}+1}$ is an explicit case of Proposition 3 with the critical sets of $c_{1}$ and $c_{2}$ being bounded, with $(q_{1},q_{2})=(0,0)$. If we consider a weaker condition $t_{0}<1$, then we can also have the Reeb digraph with exactly one edge, exactly two vertices, and the closure of the edge being homeomorphic to $D^{1}$, and in the case $t_{0}\leq 0$, the assumption (2c) of Theorem 2 (Theorem 2 (2) is also dropped.
@@
In Theorem 1, suppose that the function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function whose critical value set $c_{i}(S(c_{i}))$ is a discrete set, that at least one function of $c_{1}$ and $c_{2}$ has at least one critical point and that both the values $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ or -$\infty$ as $t$ goes to $\pm\infty$. Then the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph, where we abuse the notation. In addition, the Reeb graph is isomorphic to a graph if and only if the set $S(c_{1})\bigcup S(c_{2})\subset\mathbb{R}$ is unbounded.
@@
Case 4-1 The case $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ as $t$ goes to $\pm\infty$.
This case contains the case $c_{1}(t)$ and $c_{2}(t)$ go to -$\infty$ as $t$ goes to $\pm\infty$. In fact, it is sufficient to reverse the signs of the functions.
@@
Case 4-2 The case both $c_{1}(t)$ and $c_{2}(t)$ go to -$\infty$ ($\infty$) as $t$ goes to -$\infty$ (resp. $\infty$) .
This case contains the case $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ (-$\infty$) as $t$ goes to -$\infty$ (resp. $\infty$). In fact, it is sufficient to reverse the signs of the functions, as in Case 4-1.
@@
By rotating the two graphs in each case from Example 1 (1, 2) simultaneously and suitably, we can easily have a case for Theorem 4.
@@
In Theorem 1, suppose that the function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function the intersection of $\mathbb{R}-\{q\}$ and whose critical value set $c_{i}(S(c_{i}))$ is a discrete set and at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
    As $t\in\mathbb{R}$ goes to -$\infty$ ($\infty$), then both the values $c_{1}(t)$ and $c_{2}(t)$ converges to $q\in\mathbb{R}$.
@@
    As $t\in\mathbb{R}$ goes to $\infty$ (resp. -$\infty$), then both the values $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ or both the values go to -$\infty$.
@@
    Each connected component $K_{c_{1},c_{2},q,j}$ of the set ${{\pi}_{m+1,1}}^{-1}(q)\bigcap X_{m,c_{1},c_{2}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},q,j}$ consisting of contours of the the restriction ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and the complementary set of $K_{c_{1},c_{2},q,j}$ in the connected component $N_{c_{1},c_{2},q,j}$ consists of finitely many connected sets and contains no critical contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. In addition, these neighborhoods $N_{c_{1},c_{2},q,j}$ can be chosen to be mutually disjoint.
Then the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph. Furthermore, the Reeb graph is a graph if and only if the following hold.
@@
A proof of the former part of this theorem is shown like Theorem 2 (2) and Theorem 4.
The latter part is shown as in the proofs of several Propositions and Theorems above. As in the proofs, we discuss whether we can have a path passing an arbitrary point in an arbitrary edge of the Reeb graph connecting some two vertices in the Reeb graph and having a consistent orientation induced from the Reeb digraph. We omit precise arguments.
@@
We can find a non-negative smooth function $c_{0}(x):\mathbb{R}\rightarrow\mathbb{R}$ with $c_{0}(x):=0$, $c_{0}(x):=x+2\sin x$ ($x>T_{0}$ with a suitable $T_{0}>0$) and $c_{0}(x_{1})\leq c_{0}(x_{2})$ for an arbitrary pair of real numbers with $x_{1}<x_{2}\leq T_{0}$. Let $c_{1}(x):=\frac{1}{2(x^{2}+1)}({\sin}^{2}x)+c_{0}(x)$ and $c_{2}(x):=\frac{1}{x^{2}+1}+c_{0}(x)$. This is a case for Theorem 5 with the set $S_{(}c_{1})\bigcup S(c_{2})\subset\mathbb{R}$ being unbounded and $q=0$.
The following may be not so difficult. The author does not reach explicit answers yet.

### v4

@@ Abstract.
Hereafter, we also abuse some notation from Proposition 2 and its proof. We choose a point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$. Let $p_{D}:=(p_{D,1},p_{D,2})$. We consider a point $p_{D,c_{1},q}:=(p_{D,1},q_{D,1})$ with $c_{1}(q_{D,1})=p_{D,1}\leq p_{D,2}$ and we can have such a point due to our assumption that the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{1}$ as $t$ diverges to -$\infty$. The set $\{(p_{D,1},t)\mid q_{D,1}\leq t\leq p_{D,2}\}$ is in $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$. We can consider some path from $p_{D,c_{1},q}$ to another point $p_{s,c_{1},q}$ in $\{(x_{1},x_{2})\mid x_{2}<b_{{\rm m},0},c_{1}(x_{2})=x_{1}\}$ and have the path departing from $p_{D}$ obtained by joining the set $\{(p_{D,1},t)\mid q_{D,1}\leq t\leq p_{D,2}\}$, which is a one-point set or a segment, and the path in $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$ in such a way that the restriction of ${\pi}_{2,1}$ to the latter path is injective and that the point $(p_{s,c_{1},q},{(0)}_{j=1}^{m-1})\in{{\mathbb{R}}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$.
We discuss a fact obtained from the argument on the assumption (1) of Proposition 4 here. First, the assumption (2) of Proposition 4 is essentially same as (1) of Proposition 4. We remember some arguments in the proof of Proposition 2 (and Proposition 3) in the following arguments. From the assumptions (1, 2) of the present theorem, for a (more) general point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\overline{{D}_{m,c_{1},c_{2}}}$, we can have an oriented path from $p_{e}$ to a vertex of the Reeb graph the orientation at each point of which is same as the orientation induced from the Reeb digraph in the following way. We can choose finitely many sets of the form $\{(t,{p_{D,2}}^{\prime})\mid{p_{D,1}}^{\prime}\leq t\leq c_{2}({p_{D,2}}^{\prime})\}$ containing a point ${p_{D}}^{\prime}:=({p_{D,1}}^{\prime},{p_{D,2}}^{\prime})$ or $\{({p_{D,1}}^{\prime},t)\mid{q_{D,1}}^{\prime}\leq t\leq{p_{D,2}}^{\prime}\}$ containing a point ${p_{D}}^{\prime}:=({p_{D,1}}^{\prime},{p_{D,2}}^{\prime})$ with the relation ${p_{D,1}}^{\prime}:=c_{1}({q_{D,1}}^{\prime})$ in $\overline{{D}_{m,c_{1},c_{2}}}$ and a curve in the graph $\{(c_{i}(x),x)\mid x\in\mathbb{R}\}$ entering a point $p_{s,c}$ making $(p_{s,c},{(0)}_{j=1}^{m-1})\in{\mathbb{R}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and as a result we can have a path in $\overline{{D}_{m,c_{1},c_{2}}}$ mapped onto a desired path connecting $p_{e}$ and a desired vertex in the Reeb graph by the quotient map $q_{{\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}}:\overline{D_{c_{1},c_{2}}}\rightarrow R_{{\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}}$. The assumptions (3, 4) of the present theorem are regarded to be a kind of duals to the assumptions (1, 2) of the present theorem. We can argue as in the proofs of Propositions 2 and 3 to complete the proof of the sufficiency.
@@
Theorem 3 is also regarded as an extension of Proposition 2. Propositions 3 and 4 with their proofs are important in Theorem 3 and its proof. We only show an example for Proposition 4.
@@
The case $c_{1}(x):=\frac{t_{0}}{x^{2}+1}$ ($0<t_{0}<1$) and $c_{2}(x):=\frac{1}{x^{2}+1}$ is an explicit case of Proposition 4 with the critical sets of $c_{1}$ and $c_{2}$ being bounded and one-point sets and $(q_{1},q_{2})=(0,0)$. If we consider a weaker condition $t_{0}<1$, then we can also have the Reeb digraph with exactly one edge, exactly two vertices, and the closure of the edge being homeomorphic to $D^{1}$, and in the case $t_{0}\leq 0$, the assumption (2c) of Theorem 2 (Theorem 2 (2)) is also dropped.
@@
In Theorem 1, suppose that the set $c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2}))$ is a discrete and closed subset of $\mathbb{R}$, that at least one function of $c_{1}$ and $c_{2}$ has at least one critical point and that both the values $c_{1}(t)$ and $c_{2}(t)$ diverge to $\infty$ or -$\infty$ as $t$ diverges to $\pm\infty$. Then the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph, where we abuse the notation. In addition, the Reeb graph is isomorphic to a graph if and only if the set $S(c_{1})\bigcup S(c_{2})\subset\mathbb{R}$ is unbounded.
@@
Case 4-1 The case $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ as $t$ diverges to $\pm\infty$.
This case contains the case $c_{1}(t)$ and $c_{2}(t)$ go to -$\infty$ as $t$ diverges to $\pm\infty$. In fact, it is sufficient to reverse the signs of the functions.
@@
Case 4-2 The case both $c_{1}(t)$ and $c_{2}(t)$ go to -$\infty$ ($\infty$) as $t$ diverges to -$\infty$ (resp. $\infty$) .
This case contains the case $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ (-$\infty$) as $t$ diverges to -$\infty$ (resp. $\infty$). In fact, it is sufficient to reverse the signs of the functions, as in Case 4-1.
@@
We review Example 1. We have the first derivative of $c_{1}$ as ${c_{1}}^{\prime}(x)=\frac{2x}{{(x^{2}+1)}^{2}}{(\sin x)}^{2}+\frac{2\sin x\cos x}{x^{2}+1}$, which is bounded and converges to 0 at each infinity. We have the first derivative of $c_{2}$ as ${c_{2}}^{\prime}(x)=\frac{2x}{{(x^{2}+1)}^{2}}$, which is bounded and converges to 0 at each infinity. By rotating the two graphs in each case from Example 1 (1, 2) simultaneously, slightly, and suitably, we can easily have a case for Theorem 4 with the following conditions.
  • The resulting Reeb graph is not isomorphic to any graph.
  • The resulting case is for Case 4-2 in (the proof of) Theorem 4.
@@
In Theorem 1, suppose that the intersection $(\mathbb{R}-\{q\})\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$ is a discrete and closed subset of $\mathbb{R}-\{q\}$ and that at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
    As $t\in\mathbb{R}$ diverges to -$\infty$ ($\infty$), then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q\in\mathbb{R}$.
@@
    As $t\in\mathbb{R}$ diverges to $\infty$ (resp. -$\infty$), then both the values $c_{1}(t)$ and $c_{2}(t)$ go to $\infty$ or both the values go to -$\infty$.
@@
    Each connected component $K_{c_{1},c_{2},q,j}$ of the set ${{\pi}_{m+1,1}}^{-1}(q)\bigcap X_{m,c_{1},c_{2}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},q,j}$ in $X_{m,c_{1},c_{2}}$ consisting of contours of the the restriction ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and exactly one contour $K_{c_{1},c_{2},q,j}$ in a level set of the form ${{\pi}_{m+1,1}}^{-1}(q)\bigcap X_{m,c_{1},c_{2}}$. The complementary set of $K_{c_{1},c_{2},q,j}$ in the connected component $N_{c_{1},c_{2},q,j}$ consists of finitely many connected sets and contains no critical contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. In addition, these neighborhoods $N_{c_{1},c_{2},q,j}$ can be chosen to be mutually disjoint.
Then the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph. Furthermore, the Reeb graph is a graph if and only if the following two hold.
@@
A proof of the former part of this theorem is shown as in the proof of Theorem 2 (2) and Theorem 4.
The latter part is shown as in the proofs of several Propositions and Theorems above. As in these proofs, we discuss whether for an arbitrary point $p_{e}$ in an arbitrary edge $e$ of the Reeb graph, we can have a path passing the point $p_{e}$, containing $e$, connecting some two vertices in the Reeb graph, and having a consistent orientation induced from the Reeb digraph.
@@
We can find a non-negative smooth function $c_{0}(x):\mathbb{R}\rightarrow\mathbb{R}$ with $c_{0}(x):=0$ ($x\leq 0$), $c_{0}(x):=x+2\sin x$ ($x>T_{0}$ with a suitable $T_{0}>0$) and $c_{0}(x_{1})<c_{0}(x_{2})$ for an arbitrary pair of real numbers with $0\leq x_{1}<x_{2}\leq T_{0}$. Let $c_{1}(x):=\frac{1}{2(x^{2}+1)}({\sin}^{2}x)+c_{0}(x)$ and $c_{2}(x):=\frac{1}{x^{2}+1}+c_{0}(x)$.
This is a case for Theorem 5 with the set $S(c_{1})\bigcup S(c_{2})\subset\mathbb{R}$ being unbounded and $q=0$. This is also a case for Theorem 5 satisfying the assumptions (4, 5).
The following may not be so difficult. The author does not reach explicit answers yet.
@@
Problem 3.
Find further examples for cases of Propositions and Theorems we present here. In several cases, examples are not given in the present paper. Can we find them in the real analytic situations or using single elementary functions?